OpenStudy (anonymous):

As a punishment for something naughty that we did, my little brother and I have to whitewash both sides of a fence. We start at the same time, and we each work at a constant rate. If we each whitewash one side, I'll finish in $2$ hours and my brother will finish in $3$ hours. But I'm a nice kid, so after I finish my side, I go around to the other side and help my brother finish his side. From the time I start helping him, how many minutes does it take us to finish the job? (Just ANSWER)

OpenStudy (snuggielad):

We can't just answer the question. It's against the Code of Conduct that we all agree to upon signing up. It is punishable by banning. However, I can assist you but I need you to work with me. Where do you think you should start?

OpenStudy (anonymous):

Ok assigning variables?

OpenStudy (snuggielad):

That should work, however the dollar signs are throwing me off. Are we talking about a length is time, or price per hour?

OpenStudy (anonymous):

length of time

OpenStudy (snuggielad):

Alright, where should the variables go?

OpenStudy (anonymous):

you would use 2x and 3x

OpenStudy (anonymous):

Then...........................

OpenStudy (anonymous):

so 2x + 3x = x (I'm pretty sure)

OpenStudy (anonymous):

that doesn't make sense srry

OpenStudy (phi):

I use the idea of rate * time = 1 job and rate is in terms of "jobs per hour" and time is in hours

OpenStudy (phi):

In this problem I would call the job "paint 1 side" and the rates are \[ my\ rate = \frac{1 \ side}{2 \ hours} \\ brothers \ rate = \frac{1 \ side}{3 \ hours} \] after I finish my side. when does that happen? You should expect 2 hours, but using the equation \[ rate \cdot time = 1 \ side\\ \frac{1}{2}\cdot t = 1 \\ t= 2 \ hours\] just to show how to do it that way. Meanwhile, the other guy did some work. How much did he do in 2 hours? Using his rate we get \[ \frac{1}{3} \cdot 2 = \frac{2}{3} \ side\] He painted ⅔ of a side and there is ⅓ of a side go go.

OpenStudy (phi):

working together, they have to finish ⅓ of a side \[ \frac{1}{2} t + \frac{1}{3} t = \frac{1}{3} \] Can you finish to find t in hours? From the time I start helping him, how many minutes does it take us to finish the job? then change that into minutes